An electric field is generated in between two parallel hundred-square centimeter electrodes that are separated by a distance of two centimeters. We have one plate charged to 5.0 nano-coulombs and the other one exactly equal but opposite. An additional 'square-shaped' surface (with side length of a centimeter) is placed in between the plates and rotated such that it makes an angel of pi over 4 radians with the electric field. Determine the flux of the electric field through the square surface.
1
Expert's answer
2021-08-17T10:29:21-0400

"\Phi=ES'\cos \theta,"

"E=\frac Ud=\frac q{Cd}=\frac{qd}{d\varepsilon_0 S}=\frac q{\varepsilon_0 S},"

"\Phi=\frac{qS'}{\varepsilon_0 S}\cos \theta=8~\text{V}\cdot \text{m}."